# Security Vulnerabilities in Ethereum Smart Contracts: A Systematic Analysis

### v3

@@ 3.4.1   KotET security incident
KotET is a blockchain game. In this game, there is a “throne”, and players can send ether to the smart contract so that they can run for the throne. Assuming that player A sends 10 ETH and gets the “throne,” and then B bids 20 ETH, B inherits the “throne” and returns the 10 ETH to A. However, between February 6 and 8, 2016, players discovered a problem with the game. However, between February 6 and 8, 2016, players realized that the game had a problem, and no matter how much ETH they sent, they could not get the “throne”. This security incident was related to a DOS attack.
@@ 3.4.2   Principle of Denial of Service Attack
Denial of Service attack (Denial Of Service) means that the normal service request needed by the user cannot be processed by the system normally(De Neira et al., 2023), and the attacker launches the denial of service attack through a series of operations such as disruption, termination, freezing, etc., which leads to the normal logic of the contract can not be operated. In a smart contract, the attacker makes the user exit the contract briefly by consuming the resources of the contract, and in severe cases, it may even lead to a permanent exit, which will lock the Ethereum in the contract. There are three types of denial-of-service attacks. The first one is external manipulation or array loops, this is because mapping or array loops can be controlled externally by outsiders without a length limit, which can lead to a massive consumption of Ether and Gas, which can eventually affect the smart contract and make it suspended or permanently inoperable. The second is the state of progress of external calls, this is because the state of the smart contract is related to the execution results of external functions, but there is no protection against failure situations, which leads to the possibility of denial-of-service attacks in the event of failure of external function calls or rejection for external reasons. The third type is the owner operation, in the token contract, there will be a contract owner’s account, which has the right to suspend or execute the transaction, but if the address of the account is lost, it will lead to the whole token contract can not be executed, which causes non-subjective denial-of-service attacks.
3.5. Access Control
@@ 3.5.1    Parity Multi-Signature Wallet Theft Incident
Parity is one of the most used Ether wallets today. However, on July 19, 2017, Parity lost Ether due to a security breach (Ghaleb et al., 2023). According to the company’s report, 153,000 ETH (worth about \$30 million at the time) was confirmed stolen. On November 7, the attacker posted a question on Parity’s project website, “Anyone can kill your contract”, stating that he was not the contract owner. He said that he was not originally the owner of the contract, but the existence of the vulnerability allowed him to make himself the owner and called the contract’s kill method to erase all the code in the contract. The problem with the contract is the initialed function, which can change the owner of the contract, but there is no protection for this function, as shown in Listing 3. Also, the delegatecall() function is used in the code to make all public functions visible to anyone, as shown in Listing 4. This event is related to contractual access control.
@@ 3.5.2   Access Control Principles
When programming in Solidity, public, private, and external are the most common keywords for accessing variables or functions in Solidity. However, their roles are very different. Compare and contrast public with private: in a contract, public is the visible state of the external, while private is the visible variable of the contract instance. Access control is related to the invocation of the contract (De Oliveira et al., 2022), as shown in Figure 6.
Figure 6. Different Calling Methods
@@
3.6. Short Address Attacks
@@ 3.6.1   Principle of Short Address Attack
The ERC-20 contract will have the transfer this transaction function, its statement in the form of a function transfer(address addr,uint256 value) returns (bool success). The address is the destination address, value is the amount to be sent. As in transfer, the amount of money in the destination address of the back, so when the destination address is 0x5B38Da6a701c568545dCfcB03FcB875f56bedd00, if we will be the destination address of the last two zeros to the intentional removal the EVM is not aware of, it will be considered value in the high position in the two zeros are added. The two zeros in value are the last two zeros of addr. At this time, there will be a situation where EVM will think that we have not entered the value correctly, so it will take the liberty of adding two zeros to the back of the value, which ultimately leads to the correct address, but the amount sent has been doubled 256 times.
3.7. Advance Trading Attacks
@@ 3.7.1   Principle of Advanced Trade Attack
An early transaction attack is actually that the attacker can obtain the specific transaction information of the trader before the trader completes the operation, and then use the means of raising the offer to complete the transaction before the trader (Wu et al., 2023). This is because in Ether, all transactions have to be confirmed before they can be recorded on the chain, and each transaction carries a fee associated with it. The priority of the miner in confirming a transaction is by the level of the fee, and the higher the fee, the more the miner will deal with the transaction first. When the transaction to be confirmed is broadcast to the network, the attacker can view the details of the transaction to get the parameters they want. The process is shown in Figure 7.
Figure 7. Advance Trading Attack Process.
3.8. Privileged Function Exposure
@@ 3.8.1   Principle of Privileged Function Exposure
In Solidity, as long as the function is not modified by the privilege modifier, it will be called by everyone by default; it has a public function. The self-destruct () function (destructor) function can destroy the contract. If both of these occur in the same function, then everyone can call this function to destroy the contract. In addition to the self-destruct function, any function related to sensitive operations can also have serious consequences if exposed.
4. STRATEGY
4.1. Integer Overflow
@@ 4.1.1   Replication of integer overflow scenarios
In Listing 5, addresses a1 and a2 are the destination addresses for the transfer(Li et al., 2023d), and value1 and value2 are the amounts to be transferred from the corresponding addresses. In this code, msg. The sender is the initiator address, and require(balanceof[msg.sender]¿=value1+value2) is a judgment statement that serves to determine whether the initiator’s balance is sufficient, but there is an obvious problem with this statement. Since the upper limit of uint8 is 255, when value1 + value2¿255, the value of value1 + value2 becomes value1+value2-255. Here I set the initiator to 160 and transferred 160 and 150 to a1 and a2, respectively, and since the two add up to 310, which is more than 255. So a1+a2-256=54, in fact, the initiator only consumed 54 amount of money to a1 and a2 successfully transferred a total of 310 amount.
@@ 4.1.2   Integer Overflow Attack Solution
The most common approach is to directly call the SafeMath function(Huang et al., 2024), which is a set of smart contract functions maintained by OpenZeppelin, in addition to requiring $\text{balanceOf}[msg.sender]\geq value1+value2$ before adding a require(value1 + value 2 ¡= 255) to prevent integer overflow. 255) prevents integer overflow. Integer overflow is actually a simple arithmetic operation that has a problem, but in addition, some details of the operation are also worth noting, otherwise, it will affect the flow of the code. For example, for the array length of 256 for the unsigned integer, it is necessary to use array length++ and array length– for integer overflow check; loop variable for (var i = 1; I ¡=) items.length;i++), I is 8 for the unsigned integer, when items¿256, may lead to an I value overflow and can not be traversed completely.
4.2. Reentry Vulnerability
@@ 4.2.1   Replication of Reentry Vulnerability
In Listing 6, Balances in the original contract are used to keep track of balances, the donate function is to donate Ether to an address, the balances function is used to query how much balance the function has, and the withdraw function is used to take out a fixed amount of Ether(Fei et al., 2023). In Listing 7, Attacking the contract in Reentrancy Apple is to get an instance of the original contract, this instance is created by passing in an address, i.e., apple = Reentrancy(add). Here, the other accounts are unaware that the original contract is faulty. The attacker will first get the original contract to donate to the contract he wrote, and then, when the other accounts donate to the original contract, it is the attacker who gets the money and not the original contract account.
@@ 4.2.2   Reentry Vulnerability Solution
  • In msg.sender.call of the original contract, replace call with transfer, only then can we effectively limit the value of Gas(Tonko et al., 2024).
@@
4.3. Denial of Service Attack
@@ 4.3.1   Replication of the Denial of Service Attack
In Listing 8, we can see that under normal circumstances, if a new account enters more Ethereum (money) than the original money, it becomes the new boss, and the deposits in the original contract are returned to the previous boss. The TRANSFER transfer function is used here, which may seem safe(Mihoub et al., 2022), but it has a major problem. Because Ethernet has contract accounts as well as external accounts if the combos function calls a contract account and eventually becomes the boss, and its rollback function performs a malicious operation, the other accounts will be affected, and even if the money out is higher, they will not be able to become the new boss through the combos function.
@@ 4.3.2   Denial of Service Attack Solutions
  • For the first type of attack, the contract should adopt the retrieval mode, for example, it can adopt the function withdrawFunds(), so that it can retrieve its tokens(Uddin et al., 2024).
@@ 4.4.1   Replication of Access Control
In the contract in Listing 9, the rollback function makes a delegate call () to the control contract instance using msg. The sender is in the statement following the control double slash. In this msg. The Sender is controllable(Yang et al., 2022), and the attacker replaces msg.sender with bytes4(keccak256(“attack()”)) at this point, which allows the delegate call () call to replace the deployed control in which the owner is replaced with the attacker himself, which is msg.sender.
@@ 4.4.2   Access Control Solutions
Firstly, the delegate call () function needs to be used carefully; secondly, the control of permissions should be strengthened, and modifiers should be set for sensitive functions, such as onlyOwner(Li et al., 2023b); and lastly, to prevent the internal function from calling the external function, the internal function should also be added.
4.5. Short Address Attack
@@ 4.5.1   Short Address Attack Code Analysis
In this code in Listing 10, the approve function is a recharge function, the first parameter is the recharge address, the second parameter is the recharge amount(Kushwaha et al., 2022b), and the value of this part can be decided by yourself. The transfer is a transfer function, assuming that we call the transfer function transfer( 0x5B38Da6a701c568545dCfcB03FcB875f56bed dC4, 200). What you see inside the EVM is 0xa9059cbb00000000000000000000 32 bytes for the transfer destination address and the last 32 bytes for the transfer amount. Because in the transfer, the amount of money in the destination address the back, and is closely followed. Assuming we want to transfer the destination address 0x5B38D at this time a6a701c568545dCfcB03FcB875f56bedd00, transfer the amount of 0x1 (hexadecimal), if the destination address of the last two 0’s to remove, then the EVM will autonomously transfer the amount of the first two 0’s to the address, and then at the end of the two 0’s, at this time the transfer address is still the original address, but the amount of money transferred is not the original amount, changed to 0x5B38D. The original amount, changed to 0x100, equal to 256 times more than the original.
@@ 4.5.2   Short Address Attack Solution
4.6. Advance Trading Attacks
@@ 4.6.1   Advance Trading Attacks prevention
  • Miners can be weakened so that they no longer can sort transactions(Li et al., 2023a).
@@
In addition to this, an enhanced version of the Commit/Reveal model can be used(Arulprakash and Jebakumar, 2022). The user first encrypts the transaction and puts it on the blockchain. When this transaction has been completed, that user initiates another transaction, which is used to reveal the previous transaction, and puts it on the blockchain immediately afterward, at which point everyone can see it. But the fact that no one can know what the first transaction was used for makes it impossible for an attacker to do so blindly.
4.7. Privilege Functions
@@ 4.7.1   Replication of Privileged function
In the code in Listing 11, contract destroy1 inherits contract destroy, and before the execution of the destructor function, the destroy function in the destroy1 contract can be used normally, i.e., the name, symbol, age, and so on, these fields can be displayed normally. At this time, after calling the destructor function with another account, it destroys the destoy1 contract, which results in these fields not being displayed(Fang et al., 2023).
@@ 4.7.2   Privileged Function Preventive Measures
5. DISCUSSION AND CONCLUSION
Blockchain is a popular technology nowadays, all interested scholars want to create more value with this technology, not limited to the current field. Smart contracts based on blockchain technology also play an important role more and more, and their security or not will have a direct impact on the economy, so it is urgent to analyze the smart contract codes that have been problematic one by one(Qi et al., 2024), to understand in depth the reasons for the emergence of these vulnerabilities and to give solutions to guarantee the security of smart contracts in the future.
In this paper, we have studied the current research literature at home and abroad and summarized the security vulnerabilities of smart contracts in recent years. Utilizing the remix online compilation platform and Solidity language, this paper mainly explains the principles of integer overflow attack, re-entry attack(Wang et al., 2024), denial of service attack, and access control attack, and introduces the security events caused by them, and according to the principles of the scenarios of these vulnerabilities are reproduced, and then gives the preventive measures. In addition, short address attacks, advanced transaction attacks, and privileged function feature attacks are also introduced.
@@ 5. DISCUSSION AND CONCLUSION
Due to the promotion of blockchain technology, smart contract technology has become a hotspot of research in the industry(Li et al., 2022). However, because of the limitations of blockchain technology itself, the current smart contracts cannot meet the demands of complex logic, high-throughput data, and the lack of privacy protection for users, making their application across chains still difficult. Therefore, current smart contracts are facing the following challenges: privacy, performance, mechanism design and security, and formalized authentication. Given that smart contracts have legal issues such as unforeseen circumstances, difficulty in accountability, and lack of after-the-fact remediation, smart contracts will complement and progress in concert with traditional contracts for a long period. For smart contracts, to comprehensively maintain their legal effect(Taherdoost, 2023), smart contracts will gradually deepen the understanding of relevant laws and regulations, set standards for the review and translation of language in smart contracts, reduce translation errors, and develop standardized contract law review standards; for traditional contracts, smart contracts bring new legal application scenarios, and if you want to adapt to this new environment, you need to make the current contracts Adjustment and supplementation.
However, in addition to the problems carried out by smart contracts themselves, they also face serious challenges in the Ethereum platform. Because the debugging process of smart contracts is very complicated in the Ethereum environment(Lin et al., 2022), it is difficult to find errors in transactions. Secondly, Ether requires manual operations to establish connections with the rest of the world, which puts a lot of pressure on the developers(Li et al., 2024e). In addition to this, to set up the base measures for the Ether nodes, developers can only do it individually, which can be potentially risky. Ether is a decentralized platform, so its coding cannot be altered, and smart contract updates become a pressing issue.

### v4

@@ 3.4.1 KotET security incident
KotET is a blockchain game. In this game, there is a “throne”, and players can send ether to the smart contract so that they can run for the throne. Assuming that player A sends 10 ETH and gets the “throne,” and then B bids 20 ETH, B inherits the “throne” and returns the 10 ETH to A[44]. However, between February 6 and 8, 2016, players discovered a problem with the game. However, between February 6 and 8, 2016, players realized that the game had a problem, and no matter how much ETH they sent, they could not get the “throne”. This security incident was related to a DOS attack.
@@ 3.4.2 Principle of Denial of Service Attack
Denial of Service attack (Denial Of Service) means[46] that the normal service request needed by the user cannot be processed by the system normally[11], and the attacker launches the denial of service attack through a series of operations such as disruption, termination, freezing, etc., which leads to the normal logic of the contract can not be operated. In a smart contract, the attacker makes the user exit the contract briefly by consuming the resources of the contract, and in severe cases, it may even lead to a permanent exit, which will lock the Ethereum in the contract. There are three types of denial-of-service attacks. The first one is external manipulation or array loops, this is because mapping or array loops can be controlled externally[47] by outsiders without a length limit, which can lead to a massive consumption of Ether and Gas, which can eventually affect the smart contract and make it suspended or permanently inoperable. The second is the state of progress of external calls, this is because the state of the smart contract is related to the execution results of external functions, but there is no protection against failure situations, which leads to the possibility of denial-of-service attacks in the event of failure of external function calls or rejection for external reasons. The third type is the owner operation, in the token contract, there will be a contract owner’s account, which has the right to suspend or execute the transaction[33], but if the address of the account is lost, it will lead to the whole token contract can not be executed, which causes non-subjective denial-of-service attacks.
3.5 Access Control
@@ 3.5.1 Parity Multi-Signature Wallet Theft Incident
Parity is one of the most used Ether wallets today. However, on July 19, 2017, Parity lost Ether due to a security breach [23]. According to the company’s report, 153,000 ETH (worth about \$30 million at the time) was confirmed stolen. On November 7, the attacker posted a question on Parity’s project website, “Anyone can kill your contract”, stating that he was not the contract owner[48]. He said that he was not originally the owner of the contract, but the existence of the vulnerability allowed him to make himself the owner and called the contract’s kill method to erase all the code in the contract[49]. The problem with the contract is the initialed function[50], which can change the owner of the contract, but there is no protection for this function, as shown in Listing 3. Also, the delegatecall() function is used in the code to make all public functions visible to anyone, as shown in Listing 4. This event is related to contractual access control[50].
@@ 3.5.2 Access Control Principles
When programming in Solidity, public[52], private, and external are the most common keywords for accessing variables or functions in Solidity. However, their roles are very different. Compare and contrast public with private: in a contract, public is the visible state of the external[55], while private is the visible variable of the contract instance. Access control is related to the invocation of the contract [12], as shown in Figure 6.
Figure 6: Different Calling Methods
@@
3.6 Short Address Attacks
@@ 3.6.1 Principle of Short Address Attack
The ERC-20 contract will have the transfer this transaction function, its statement in the form of a function transfer(address addr,uint256 value) returns (bool success). The address is the destination address, value is the amount to be sent. As in transfer, the amount of money in the destination address of the back, so when the destination address is 0x5B38Da6a701c568545d
CfcB03FcB875f56bedd00, if we will be the destination address of the last two zeros to the intentional removal the EVM is not aware of, it will be considered value in the high position in the two zeros are added. The two zeros in value are the last two zeros of addr. At this time, there will be a situation where EVM will think that we have not entered the value correctly, so it will take the liberty of adding two zeros to the back of the value, which ultimately leads to the correct address, but the amount sent has been doubled 256 times.
3.7 Advance Trading Attacks
@@ 3.7.1 Principle of Advanced Trade Attack
An early transaction attack is actually that the attacker can obtain the specific transaction information of the trader before the trader completes the operation[56], and then use the means of raising the offer to complete the transaction before the trader [71]. This is because in Ether, all transactions have to be confirmed before they can be recorded on the chain, and each transaction carries a fee associated with it. The priority of the miner in confirming a transaction is by the level of the fee, and the higher the fee, the more the miner will deal with the transaction first. When the transaction to be confirmed is broadcast to the network[58], the attacker can view the details of the transaction to get the parameters they want. The process is shown in Figure 7.
Figure 7: Advance Trading Attack Process.
3.8 Privileged Function Exposure
@@ 3.8.1 Principle of Privileged Function Exposure
In Solidity, as long as the function is not modified by the privilege modifier, it will be called by everyone by default; it has a public function. The self-destruct () function (destructor) function can destroy the contract[57]. If both of these occur in the same function, then everyone can call this function to destroy the contract. In addition to the self-destruct function, any function related to sensitive operations can also have serious consequences if exposed.
4 STRATEGY
4.1 Integer Overflow
@@ 4.1.1 Replication of integer overflow scenarios
In Listing 5, addresses a1 and a2 are the destination addresses for the transfer[37], and value1 and value2 are the amounts to be transferred from the corresponding addresses. In this code, msg. The sender is the initiator address, and require(balanceof[msg.sender]¿=value1+value2) is a judgment statement that serves to determine whether the initiator’s balance is sufficient, but there is an obvious problem with this statement. Since the upper limit of uint8 is 255, when value1 + value2¿255, the value of value1 + value2 becomes value1+value2-255. Here I set the initiator to 160 and transferred 160 and 150 to a1 and a2, respectively, and since the two add up to 310, which is more than 255. So a1+a2-256=54, in fact, the initiator only consumed 54 amount of money to a1 and a2 successfully transferred a total of 310 amount.
@@ 4.1.2 Integer Overflow Attack Solution
The most common approach is to directly call the SafeMath function[28], which is a set of smart contract functions maintained by OpenZeppelin, in addition to requiring $\text{balanceOf}[msg.sender]\geq value1+value2$ before adding a require(value1 + value 2 ¡= 255) to prevent integer overflow. 255) prevents integer overflow. Integer overflow is actually a simple arithmetic operation that has a problem, but in addition, some details of the operation are also worth noting, otherwise, it will affect the flow of the code. For example, for the array length of 256 for the unsigned integer, it is necessary to use array length++ and array length– for integer overflow check; loop variable for (var i = 1; I ¡=) items.length;i++), I is 8 for the unsigned integer, when items¿256, may lead to an I value overflow and can not be traversed completely.
4.2 Reentry Vulnerability
@@ 4.2.1 Replication of Reentry Vulnerability
In Listing 6, Balances in the original contract are used to keep track of balances[60], the donate function is to donate Ether to an address, the balances function is used to query how much balance the function has, and the withdraw function is used to take out a fixed amount of Ether[18]. In Listing 7, Attacking the contract in Reentrancy Apple is to get an instance of the original contract, this instance is created by passing in an address, i.e., apple = Reentrancy(add). Here, the other accounts are unaware that the original contract is faulty. The attacker will first get the original contract to donate to the contract he wrote, and then, when the other accounts donate to the original contract, it is the attacker who gets the money and not the original contract account[62].
@@ 4.2.2 Reentry Vulnerability Solution
  • 1.
    In msg.sender.call of the original contract, replace call with transfer, only then can we effectively limit the value of Gas[67].
  • 2.
@@
4.3 Denial of Service Attack
@@ 4.3.1 Replication of the Denial of Service Attack
In Listing 8, we can see that under normal circumstances[65], if a new account enters more Ethereum (money) than the original money, it becomes the new boss, and the deposits in the original contract are returned to the previous boss. The TRANSFER transfer function is used here, which may seem safe[51], but it has a major problem. Because Ethernet has contract accounts as well as external accounts if the combos function calls a contract account and eventually becomes the boss, and its rollback function performs a malicious operation, the other accounts will be affected, and even if the money out is higher, they will not be able to become the new boss through the combos function[64].
@@ 4.3.2 Denial of Service Attack Solutions
  • 1.
    For the first type of attack, the contract should adopt the retrieval mode, for example, it can adopt the function withdrawFunds(), so that it can retrieve its tokens[68].
  • 2.
@@ 4.4.1 Replication of Access Control
In the contract in Listing 9, the rollback function makes a delegate call () to the control contract instance using msg. The sender is in the statement following the control double slash. In this msg. The Sender is controllable[72], and the attacker replaces msg.sender with bytes4(keccak256(“attack()”)) at this point, which allows the delegate call () call to replace the deployed control in which the owner is replaced with the attacker himself, which is msg.sender.
@@ 4.4.2 Access Control Solutions
Firstly, the delegate call () function needs to be used carefully; secondly, the control of permissions should be strengthened, and modifiers should be set for sensitive functions, such as onlyOwner[39]; and lastly, to prevent the internal function from calling the external function, the internal function should also be added.
4.5 Short Address Attack
@@ 4.5.1 Short Address Attack Code Analysis
In this code in Listing 10, the approve function is a recharge function, the first parameter is the recharge address, the second parameter is the recharge amount[36], and the value of this part can be decided by yourself. The transfer is a transfer function, assuming that we call the transfer function transfer( 0x5B38Da6a701c568545dCfcB03FcB875f56bed dC4, 200). What you see inside the EVM is 0xa9059cbb00000000000000000000 32 bytes for the transfer destination address and the last 32 bytes for the transfer amount. Because in the transfer, the amount of money in the destination address the back, and is closely followed. Assuming we want to transfer the destination address 0x5B38D at this time a6a701c568545dCfcB03FcB875f56bedd00, transfer the amount of 0x1 (hexadecimal), if the destination address of the last two 0’s to remove, then the EVM will autonomously transfer the amount of the first two 0’s to the address, and then at the end of the two 0’s, at this time the transfer address is still the original address, but the amount of money transferred is not the original amount, changed to 0x5B38D. The original amount, changed to 0x100, equal to 256 times more than the original.
@@ 4.5.2 Short Address Attack Solution
4.6 Advance Trading Attacks
@@ 4.6.1 Advance Trading Attacks prevention
  • 1.
    Miners can be weakened so that they no longer can sort transactions[38].
  • 2.
@@
In addition to this, an enhanced version of the Commit/Reveal model can be used[2]. The user first encrypts the transaction and puts it on the blockchain. When this transaction has been completed, that user initiates another transaction, which is used to reveal the previous transaction, and puts it on the blockchain immediately afterward, at which point everyone can see it. But the fact that no one can know what the first transaction was used for makes it impossible for an attacker to do so blindly.
4.7 Privilege Functions
@@ 4.7.1 Replication of Privileged function
In the code in Listing 11, contract destroy1 inherits contract destroy, and before the execution of the destructor function, the destroy function in the destroy1 contract can be used normally, i.e., the name, symbol, age, and so on, these fields can be displayed normally. At this time, after calling the destructor function with another account, it destroys the destoy1 contract, which results in these fields not being displayed[17].
@@ 4.7.2 Privileged Function Preventive Measures
5 DISCUSSION AND CONCLUSION
Blockchain is a popular technology nowadays, all interested scholars want to create more value with this technology, not limited to the current field. Smart contracts based on blockchain technology also play an important role more and more, and their security or not will have a direct impact on the economy, so it is urgent to analyze the smart contract codes that have been problematic one by one[54], to understand in depth the reasons for the emergence of these vulnerabilities and to give solutions to guarantee the security of smart contracts in the future.
In this paper, we have studied the current research literature at home and abroad and summarized the security vulnerabilities of smart contracts in recent years. Utilizing the remix online compilation platform and Solidity language, this paper mainly explains the principles of integer overflow attack, re-entry attack, denial of service attack, and access control attack, and introduces the security events caused by them, and according to the principles of the scenarios of these vulnerabilities are reproduced, and then gives the preventive measures. In addition, short address attacks, advanced transaction attacks, and privileged function feature attacks are also introduced.
@@ 5 DISCUSSION AND CONCLUSION
Due to the promotion of blockchain technology, smart contract technology has become a hotspot of research in the industry[40]. However, because of the limitations of blockchain technology itself, the current smart contracts cannot meet the demands of complex logic, high-throughput data, and the lack of privacy protection for users, making their application across chains still difficult. Therefore, current smart contracts are facing the following challenges: privacy, performance, mechanism design and security, and formalized authentication. Given that smart contracts have legal issues such as unforeseen circumstances, difficulty in accountability, and lack of after-the-fact remediation, smart contracts will complement and progress in concert with traditional contracts for a long period. For smart contracts, to comprehensively maintain their legal effect[66], smart contracts will gradually deepen the understanding of relevant laws and regulations, set standards for the review and translation of language in smart contracts, reduce translation errors, and develop standardized contract law review standards; for traditional contracts, smart contracts bring new legal application scenarios, and if you want to adapt to this new environment, you need to make the current contracts Adjustment and supplementation.
However, in addition to the problems carried out by smart contracts themselves, they also face serious challenges in the Ethereum platform. Because the debugging process of smart contracts is very complicated in the Ethereum environment[45], it is difficult to find errors in transactions. Secondly, Ether requires manual operations to establish connections with the rest of the world, which puts a lot of pressure on the developers[42]. In addition to this, to set up the base measures for the Ether nodes, developers can only do it individually, which can be potentially risky. Ether is a decentralized platform, so its coding cannot be altered, and smart contract updates become a pressing issue.